A boy Is 10 years old and his mother is 37 years old. In how many years time will the mother be twice as old as her son
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52776)

.

Let the unknown number of years will be x.


Then you have this equation

    37 + x = 2*(10+x)


Simplify and find x

    37 + x = 20 + 2x

    37 - 20 = 2x - x

       17   =    x.


ANSWER.  In 17 years.

Solved.



Answer by greenestamps(13198)



While a formal algebraic solution was probably wanted, you can get good problem-solving experience and good mental exercise by solving the problem using logical reasoning and simple arithmetic.

The difference in their ages is 37-10 = 27, so when the mother will be twice as old as her son their ages will be 54 and 27. Since the mother is now 37, that will be in 54-37 = 17 years.

ANSWER: 17 years